# ElectricityEmissions.jl: A Framework for the Comparison of Carbon Intensity Signals

### v2

@@ Appendix B LMCE Calculation
where $A\in\mathbb{R}^{n\times n}$ contains the problem’s equality and binding inequality constraints, $x^{*}\in\mathbb{R}^{n}$ contains the decision variable values at optimality, and $b\in\mathbb{R}^{n}$ contains the standard-form RHS constraint bounds.
@@
where $\Theta\in\mathbb{R}^{N}$ refers to the vector of bus voltage angles, $P_{G}\in\mathbb{R}^{N_{G}}$ the vector of generator set-points, and $C_{g}\in\mathbb{R}^{N_{G}}$ the generator cost values. $P_{D}\in\mathbb{R}^{N}$ is the vector containing total load at each bus.
In this formulation, the nodal power balance constraints make up the first $N$ rows of the $A$ matrix (the only constraints to include $P_{D}$ variables). For ease in calculation, they are ordered according to their bus index. The final $n-N$ rows of $A$ correspond to the slack bus equality constraints, and all binding inequality constraints (for which the order does not matter).
We then invert the $A$ matrix, and solve for $\Delta P_{D}$. For compactness, we define the $B$ matrix to be the first $N$ columns and rows $N+1$ through $N+N_{g}$ of $A^{-1}$ These are simply the columns of $A^{-1}$ aligning with $\Delta P_{D}$ and the rows aligning with $\Delta P_{G}$ We can then write:
(10) $\Delta P_{G}=B\Delta P_{D}$
To determine the change in generation resulting from one unit of additional load at bus $i$ ($\Delta p_{G,i}$), we can set $\Delta P_{D}$ to be the ”one-hot” encoding of the bus in question (setting the $i$th entry to 1 and all others to 0). We can then perform the final LMCE calculation:
(11) $LMCE_{i}=\langle\Delta p_{G,i},e_{g}\rangle$
[…]
Appendix C Load Shifting Formulation

### v3

@@ A.2. LMCE Calculation
where $A\in\mathbb{R}^{n\times n}$ contains the problem’s equality and binding inequality constraints, $x^{*}\in\mathbb{R}^{n}$ contains the (unique) decision variable values at optimality, and $b\in\mathbb{R}^{n}$ contains the standard-form RHS constraint bounds.
We note that not all solutions returned by a solver may be basic feasible, and it is possible that the problem exhibits degeneracy. One notable potential source of degeneracy is a non-unique solution due to the presence of more than one generator with the same cost function at a given node, leading to an A matrix that is not full rank. As noted in the first step of the example ElectricityEmissions.jl workflow (Section 4.3), this is easily remedied by adding a small amount of noise to such duplicate cost functions.
@@
where $\Theta\in\mathbb{R}^{N}$ refers to the vector of bus voltage angles, $p_{G}\in\mathbb{R}^{N_{G}}$ the vector of generator set-points, and $C_{g}\in\mathbb{R}^{N_{G}}$ the generator cost values. $p_{D}\in\mathbb{R}^{N}$ is the vector containing total load at each bus. In this formulation, the nodal power balance constraints, which are the only constraints to include $P_{D}$ variables, make up the first $N$ rows of the $A$ matrix. For ease in calculation, they are ordered according to their bus index. The final $n-N$ rows of $A$ correspond to the slack bus equality constraints, and all binding inequality constraints for which the order does not matter. This system of equations is identical to that used in (Lindberg et al., 2021a), save for the addition of the $C_{G}$ variables representing the piecewise-linear generator costs, and any binding constraints on these variables (encoded in the $A$ matrix).
Following the formation of this system, we then invert the $A$ matrix, and solve for $\Delta p_{D}$. For compactness, we define the $B$ matrix to be the first $N$ columns and rows $N+1$ through $N+N_{g}$ of $A^{-1}$ These are simply the columns of $A^{-1}$ aligning with $\Delta p_{D}$ and the rows aligning with $\Delta p_{G}$ We can then write
(10) $\Delta p_{G}=B\Delta p_{D}.$
To determine the change in generation $\Delta p_{G,j}$ at generator $j$ resulting from one unit of additional load at bus $i$, we set the $i$th entry of $\Delta p_{D}$ to 1 and all others to 0. We can then perform the final LMCE calculation as
(11) $LMCE_{i}=\sum_{j\in\mathcal{G}}e_{G,j}\Delta p_{G,j},$
where $e_{G,j}$ generator emissions intensity of generator $j$.
A.3. Load Shifting Formulation
@@
We note that shifting outcomes, including those presented in this work, are likely to vary significantly depending on the specific implementation. Of particular importance are factors that influence the size of load-shifts, such as constraints on data-center flexibility ($\epsilon$ in our model), as well as case-specific shifting costs (e.g. from communications overhead) which we do not consider here. As a result of such effects, we do not include the shifting formulation employed in this work as part of the ElectricityEmissions.jl package, under the assumption that users may wish to model a variety of flexible load types with specific qualities/requirements.
Los Alamos Unlimited Release LA-UR-24-31644. Reviewed for release outside the Laboratory with no distribution restrictions.